OpenStudy (anonymous):

solve each problem by elimination -5x-y=3 -6x+y=8 and show your work

OpenStudy (espex):

Where are you stuck?

OpenStudy (anonymous):

With everything. I just need someone to explain the whole process through to me. And with numbers. Its all so complicated.

OpenStudy (espex):

What you are trying to do is "eliminate" one of the variables. You do this by adding or subtracting the two equations with respect to each other.

OpenStudy (espex):

For example, you have -5x-y=3 -6x+y=8 Adding them together ---------------- -11x +0y=11

OpenStudy (espex):

Solve for x.

OpenStudy (anonymous):

Can you help me solve x please.

OpenStudy (espex):

How would you get x alone?

OpenStudy (anonymous):

I don't know maybe you can explain that to me

OpenStudy (espex):

if you have -11x=11, what would you do to get x by itself?

OpenStudy (anonymous):

So would it be -1

OpenStudy (espex):

Exactly.

OpenStudy (anonymous):

Now can you help me with y